OpenStudy (anonymous):

1. Solve: y2 + 7x = 30 2. Solve: x2 + 7 = -8x 3. Solve: x2 – 26x + 48 = 0 4. Solve: x2 - 4x = 32 5. Solve: w(6w + 12) = 0 6. Solve: y2 + y = 30 7. Solve: (a – 5)(a + 3) = 0

Parth (parthkohli):

That's a lot to solve hehe!

Parth (parthkohli):

Do you know anything about quadratic equations?

OpenStudy (anonymous):

No not really

Parth (parthkohli):

All right. You're ready to embark on a journey :)

Parth (parthkohli):

There are infinite number of ways to solve a quadratic equation, but you may use the formula to start.

Parth (parthkohli):

A quadratic equation looks like: \(ax^2 + bx + c = 0\) It has 2 number of solutions(maximum)

OpenStudy (anonymous):

ok

Parth (parthkohli):

And the formula to solve for \(x\) is: \( \color{Black}{\Rightarrow \Large x = {-b \pm \sqrt{b^2 - 4ac} \over 2a} }\)

OpenStudy (anonymous):

right ..

Parth (parthkohli):

Do you know what \(\pm\) means?

OpenStudy (anonymous):

yess

Parth (parthkohli):

Yes, anyway, it means that x as two solutions: \( \color{Black}{\Rightarrow \Large {x = {-b + \sqrt{b^2 - 4ac} \over 2a}} }\) OR \( \color{Black}{\Rightarrow \Large x = {-b - \sqrt{b^2 - 4ac} \over 2a} }\)

OpenStudy (anonymous):

right soo then i just plug in ?

Parth (parthkohli):

yep.

OpenStudy (anonymous):

Thanks

Parth (parthkohli):

There's another thing here. The zero product rule.

OpenStudy (anonymous):

oh whats that ?

Parth (parthkohli):

Some of them are given like (a - 5)(a + 3) = 0 Zero product rule says that if a * b = 0, then either a = 0 or b = 0

Parth (parthkohli):

This is because any number multiplied by 0 is 0! :)

Parth (parthkohli):

So for example \((a - 5)(a + 3) = 0 \) Either a - 5 is 0 or a + 3 is 0. We can equate.

Parth (parthkohli):

\( \color{Black}{\Rightarrow a - 5 = 0 }\) \( \color{Black}{\Rightarrow a = 5 }\) OR \( \color{Black}{\Rightarrow a + 3 = 0 }\) \( \color{Black}{\Rightarrow a = -3 }\)

OpenStudy (anonymous):

Ohh I get it . that means the problem cant be solved ?

Parth (parthkohli):

It can be solved.

Parth (parthkohli):

That's when another method known as factorization comes in, though it may be difficult for me to explain.

Parth (parthkohli):

Did you get that (a - 5)(a + 3) = 0 example I gave you?

OpenStudy (anonymous):

Yes I get it . (:

Parth (parthkohli):

Aha! Is there anything left for me to explain?

OpenStudy (anonymous):

i think i have the hang of it . (: Thankyou SoMUCH !

Parth (parthkohli):

You're welcome :)